# Inflationary Krylov complexity

### v2

@@ V.1 Generalized Lanczos algorithm
Recalling that our Hamiton (23) is Hermitian. For the closed part (35), it only contains the two terms which are proportional to $\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}$, respectively. For us, the rest two terms which are proportional to $\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}$ and $\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}$, they are responsible for producing $c_{n}$ which will be confirmed in the latter investigation. As Socorro:2022aoz discussed, the k-essence gravitational system is non-Hermitetian, it strongly indicates the Hamilton of inflationary model is non-Hermitian. However, our studying Hamilton (23) is Hermitian, such as $v^{3}$, $\partial_{i}vv^{2}$ $\it e.t.c,$ that could give the non-Hermitian terms of Hamilton. The problem is that the action (16) is only the quadratic perturbation of Mukhanov variable $v=z\mathcal{R}$, the higher order perturbation of $v$ could give rise to the non-Hermitian terms. Meanwhile, we also neglect the potential of inflation describing the interaction between the dynamic system and the enviroment. Here we give more details for readers who are not familiar with the perturbation theory of inflation. In light of Baumann:2009ds , we have the quadratic action in terms of the Mukhanov variable as shown
@@
in momentum space where $\phi$ is the inflaton field (background field). If the potential has the he Yukawa coupling term $\lambda_{s}\phi^{2}\chi^{2}$ and we keep it in the action that also will lead to the Non-Hermitian terms. Due to the slow-roll conditions $\frac{a^{\prime\prime}}{a}\gg\frac{d^{2}V}{d\phi^{2}}$, one could neglect the contribution of potential.
@@ V.3 Wave function of $\phi(n)$ in a weak dissipative system
We do a simple analysis for $\chi\mu$. Here, $\chi$ is a free parameter. In the later investigation, we will show that $\chi\mu$ will play a role in the dissipative coefficient. For completely solving PDE (49), we will make a Fourier transformation between the conformal time and frequency,
@@ V.4 Exact wave function of $\phi(n)$ in an open inflationary universe
Before constructing the exact wave function of $\phi_{n}$. We will present the validity of the method coming via Parker:2018yvk , in which they stated that their whole analysis is in light of the asymptotic behavior of $b_{n}^{(1)}$,
@@
where $\alpha$ and $\gamma$ are real constants determined by various models. This kind of behavior of $b_{n}$ is quite universal for an infinite, non-integrable, many-body system. In our case, the $b_{n}$ (36) is of the linear growth depending on $n$ which is perfected and agreed with their assumption. Although this kind of method is for the closed system, one can naturally promote it into the open system Bhattacharjee:2022lzy due to the high symmetry of $\mathcal{L}_{o}$ corresponding to the total Hamilton (23). Based on these two points, we will use the generalized way of constructing the exact wave function. First, we will define the two parameters as follows:
@@
where $b_{n}$ and $c_{n}$ (corresponding to $a_{n}$ of Bhattacharjee:2022lzy ) are valid for many specific models, only requiring that $b_{n}\propto n$, and $u_{1}$ and $u_{2}$ encoding the information of various models. Comparing with (36) and (48), one can straightforwardly obtain
@@ V.5 Exact evolution of Krylov complexity in an open inflationary universe
Once obtaining the exact formula of Krylov complexity, one can precisely investigate the evolution of Krylov complexity. We have analyzed that the values of $u_{1}$ and $u_{2}$ will be larger than one. Observing that the values of $u_{1}$ and $u_{2}$ are mainly determined by $k$ and $a$. Our numerics reveals that the $u_{1}$ and $u_{2}$ will be smaller than one as $k<0.01$ and $-4<y<0$, which can be a criterion for assessing whether $u_{1}$ and $u_{2}$ are larger than one or not. Recalling that the universe undergoes an exponential expansion during an inflationary period, the small values of $k$ could represent the evolution of Krylov complexity as inflation almost finishes. At the very beginning of inflation, the values of $k$ should be large. Consequently, one could see a complete evolution of Krylov complexity when $k$ varies from tiny values to large values.
@@
First, we will give the precise evolution of Krylov complexity in terms of (60) represented by Fig. 4. One can see that the total trend of Krylov complexity is almost the same with various values of $\alpha$, the only difference is the order of the Krylov complexity, the standard dispersion relation will hold the largest Krylov complexity under the same parameters, which is contradicted with the evolution in closed system presenting in Fig. 3, where the modified case will grow faster compared with the standard case. Observing that there is a blank area for every case, which means that the Krylov complexity will decrease to zero and then it will experience exponential growth. One can also find that the varying trend of Krylov complexity is mainly determined by scale factor $a$, not $k$. In light of this observation, we could fix some specific values of $k$ and vary with respect to $y$ in order to check the evolution trend of Krylov complexity. For completeness, we only consider two cases: (a). $k=0.01$; (b). $k=1000$ which could provide a full perspective of the evolution trend. Fig. 5 clearly indicates that the Krylov complexity will reach the peak of various $\alpha$, then it will dramatically decrease to zero. After some time scale, it will show the exponential growth pattern, in which the standard case will grow faster compared with modified cases. Thus, Fig. 5 will confirm the varying trend as shown in Fig. 3. In some sense, Fig. 5 has given a precise and complete evolution of Krylov complexity.
@@ V.6 Krylov entropy in an open inflationary universe
We have checked that the wave function (58) could nicely recover the case of the weak region, meanwhile it also can restore into the case closed system. For completeness, we will also calculate its corresponding Krylov entropy.
Being armed with the definition of Krylov entropy (39) and its resulting formula is derived in Appendix A that can be read as follows,
@@
$\begin{split}&S_{k}\approx-\big{(}\cosh^{2}\eta-4u_{2}\cosh^{3}\sinh\eta\big{)}\bigg{[}-2u_{2}\tanh\eta+\rm sech^{2}\eta\ln\rm sech^{2}\eta+2u_{2}\tanh\eta\ln\rm sech^{2}\eta\\ &+\tanh^{2}\eta\ln\tanh^{2}\eta\bigg{]}\\ &=-\sinh^{2}\eta\ln\tanh^{2}\eta-\ln\rm sech^{2}\eta+u_{2}\big{(}2\cosh\eta\sinh\eta+2\cosh\eta\sinh\eta\ln\rm sech^{2}\eta\\ &+4\cosh\eta\sinh^{3}\eta\ln\tanh^{2}\eta\big{)}\\ &=\cosh^{2}\eta\ln\cosh^{2}\eta-\sinh^{2}\eta\ln\sinh^{2}\eta+u_{2}\big{(}2\cosh\eta\sinh\eta+2\cosh\eta\sinh\eta\ln\rm sech^{2}\eta\\ &+4\cosh\eta\sinh^{3}\eta\ln\tanh^{2}\eta\big{)},\end{split}$ (63)
@@
$(a).$ Sec. II gives a brief introduction to Krylov complexity and calculates the Krylov complexity and Lanczos coefficient in the closed system’s method. Sec. III simply introduces what is the Krylov complexity and Lanczos coefficient in a closed system, then we also give some very fundamental information about inflation whose more details can be found in Baumann:2009ds . Based on these two concepts, we found that $b_{n}$ is proportional to $n$ as shown in Eq. (36), thus it could be seen that our universe is an infinite, many-body, and chaotic system confirmed by Bhattacharya:2022gbz . Meanwhile, Fig. 2 indicates that the standard dispersion case of the Lanczos coefficient and Lyapunov index are mainly determined by scale factor (single field inflation), the modified case is mainly determined by momentum $k$, where these two parameters present the chaotic features of the system. To be more precise, the scale factor could mainly determine the growth of the operator and chaos for the standard case. As for the modified case, it was mostly influenced by the momentum.
@@
$(b).$ Fig. 3 clearly indicates that the Krylov complexity will show the irregular oscillation before the horizon exits which shows a similar pattern compared with circuit complexity Li:2023ekd . And the modified case will lead to the faster growth of complexity (exponential growth). We also found that the number density is the same with Krylov complexity similar to Adhikari:2022oxr . Since the number density is proportional to volume, we illustrate the CV conjecture in some sense. Ref. Barbon:2019wsy shows there is a saturation value of Krylov complexity. In our case, it may show a similar trend of complexity by extending our analysis into RD and MD. Meanwhile, what we consider is a closed system.
@@
$(c).$ In the approach of an open system, the investigation is based on our central result, an exact wave function (58) describing the strong and weak dissipative system, including inflation. Compared with the original method for constructing the wave function Bhattacharjee:2022lzy , our wave function (58) could describe the strong and weak dissipative region. According to $\eta=-\frac{1}{aH}$, one can see that $\eta$ cannot be larger than zero in inflation. But the situation will be different after inflation since the $\eta\propto t^{\alpha}$ in the late universe. Thus, one can conclude that the inflation is a strong dissipative system. However, the analysis can be dubbed as a criterion for testing the validity of the exact wave function (58). As a further test, the calculation for Krylov complex (61) and Krylov entropy (63) could nicely recover into the case of a closed system. In the weak dissipative region, our calculation could return into the analysis of Bhattacharjee:2022lzy , where there is scrambling time dubbed as a boundary for studying the complexity. The scrambling time is just $\xi$ (54) which is a variable determined by $k$ and $a$. For the strong dissipative region, one could see that the varying trend is always growing from the large scales, which will not reach some constant values. Fig. 4 clearly indicates the evolution trend of Krylov complexity. Generally, one cannot distinguish the standard case and modified case from the strong dissipative region. At the small scale, there is also one peak of Krylov complexity. One can see that inflation will have a huge impact on the Krylov complexity.
@@
Secondly, all of our analysis is based on the Hamilton (23) that is Hermitian. The original method for evaluating the Krylov complexity in an open system is based on the Lindblad master equation Lindblad:1975ef ; Gorini:1975nb . The corresponding Lindblad of various dynamical systems is non-Hermitian. For obtaining the non-Hermitian Hamilton, one could expand the action into the higher order in terms of the Mukhanov variable. Another way is that we could add the potential of inflaton in MD or RD. The contribution of potential cannot be ignored due to the violation of slow-roll condition.
Thirdly, the horizon exits will lead to the curvature perturbation transit from the quantum level to the classical level. Thus, this process could impact the evolution of Krylov complexity. Especially, one can expand the action of inflaton to the higher order to produce the minimal decoherence Burgess:2022nwu , which only incorporates the self-interaction via General Relativity. We are still curious about whether the holographic principle can be applied to this inflationary period or not, just like the CV or CA conjecture.
@@ Appendix A The calculation of Krylov entropy
$\begin{split}&S_{k}=-\frac{1}{A^{2}}\bigg{[}\rm sech^{2}\eta[1+2u_{2}\tanh\eta+(u_{2}^{2}-|1-u_{1}^{2}|)\tanh^{2}\eta][\ln\rm sech^{2}-\ln(1+u_{2}\tanh\eta)^{2}]\\ &+\rm sech^{2}\eta\tanh^{2}\eta|1-u_{1}^{2}|\ln|1-u_{1}^{2}|+|1-u_{1}^{2}|\tanh^{2}\eta\rm sech^{2}\eta[\ln\tanh^{2}\eta-\ln(1+u_{2}\tanh\eta)^{2}]\bigg{]}\\ &=-\frac{1}{A^{2}}\bigg{[}-\ln(1+u_{2}\tanh\eta)^{2}(\rm sech^{2}\eta+2u_{2}\rm sech^{2}\eta\tanh\eta+u_{2}^{2}\rm sech^{2}\eta\tanh^{2}\eta)\\ &+\rm sech^{2}\eta\ln\rm sech^{2}\eta[1+2u_{2}\tanh\eta+(u_{2}^{2}-|1-u_{1}^{2}|)\tanh^{2}\eta]+\rm sech^{2}\eta\tanh^{2}\eta|1-u_{1}^{2}|\ln|1-u_{1}^{2}|]\\ &+|1-u_{1}^{2}|\tanh^{2}\eta\rm sech^{2}\eta\ln\tanh^{2}\eta\bigg{]}\\ &=\frac{\rm sech^{2}\eta}{A^{2}}\bigg{[}(1+u_{2}\tanh\eta)^{2}\ln(1+u_{2}\tanh\eta)^{2}-[(1+u_{2}\tanh\eta)^{2}-|1-u_{1}^{2}|\tanh^{2}\eta]\ln\rm sech^{2}\eta\\ &-\tanh^{2}\eta|1-u_{1}^{2}|\ln|1-u_{1}^{2}|-|1-u_{1}^{2}|\tanh^{2}\eta\ln\tanh^{2}\eta\bigg{]}.\end{split}$ (70)

### v4

@@ V.1 Generalized Lanczos algorithm
Recalling that our Hamiton (23) is Hermitian. For the closed part (35), it only contains the two terms which are proportional to $\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}$, respectively. For us, the rest two terms which are proportional to $\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}$ and $\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}$, they are responsible for producing $c_{n}$ which will be confirmed in the latter investigation. As Socorro:2022aoz discussed, the k-essence gravitational system is non-Hermitetian, it strongly indicates the Hamilton of inflationary model is non-Hermitian. However, our studying Hamilton (23) is Hermitian, the problem is that the action (16) is only the quadratic perturbation of Mukhanov variable $v=z\mathcal{R}$. The higher order perturbation of $v$ could give rise to the non-Hermitian terms, such as $v^{3}$, $\partial_{i}vv^{2}$ $\it e.t.c,$ that could give the non-Hermitian terms of Hamilton. Meanwhile, we also neglect the potential of inflation describing the interaction between the dynamic system and the enviroment. Here we give more details for readers who are not familiar with the perturbation theory of inflation. In light of Baumann:2009ds , we have the quadratic action in terms of the Mukhanov variable as shown
@@
in momentum space where $\phi$ is the inflaton field (background field). If the potential has the Yukawa coupling term $\lambda_{s}\phi^{2}\chi^{2}$ and we keep it in the action that also will lead to the Non-Hermitian terms. Due to the slow-roll conditions $\frac{a^{\prime\prime}}{a}\gg\frac{d^{2}V}{d\phi^{2}}$, one could neglect the contribution of potential.
@@ V.3 Wave function of $\phi(n)$ in a weak dissipative system
We do a simple analysis for $\chi\mu$. Here, $\chi$ is a free parameter. In the later investigation, we will show that $\chi\mu$ will play a role in the dissipative coefficient. For completely solving PDE (49), we will make a Fourier transformation between $\eta$ and frequency $\omega$,
@@ V.4 Exact wave function of $\phi(n)$ in an open inflationary universe
Before constructing the exact wave function of $\phi_{n}$, we will present the validity of the method coming via Parker:2018yvk , in which they stated that their whole analysis is in light of the asymptotic behavior of $b_{n}^{(1)}$ (Lanczos coefficient),
@@
where $\alpha$ and $\gamma$ are real constants determined by various models. This kind of behavior of $b_{n}$ is quite universal for an infinite, non-integrable, many-body system. In our case, the $b_{n}$ (36) is of the linear growth depending on $n$ which is perfectly agreed with their assumption. Although this kind of method is for the closed system, one can naturally promote it into the open system Bhattacharjee:2022lzy due to the high symmetry of $\mathcal{L}_{o}$ corresponding to the total Hamilton (23). Based on these two points, we will use the generalized way of constructing the exact wave function. First, we will define the two parameters as follows:
@@
where $b_{n}$ and $c_{n}$ (corresponding to $a_{n}$ of Bhattacharjee:2022lzy ) are valid for many specific models, only requiring that $b_{n}\propto n$, and $u_{1}$ and $u_{2}$ are encoding the information of various models. Comparing with (36) and (48), one can straightforwardly obtain
@@ V.5 Exact evolution of Krylov complexity in an open inflationary universe
Once obtaining the exact formula of Krylov complexity, one can precisely investigate the evolution of Krylov complexity. We have analyzed that the values of $u_{1}$ and $u_{2}$ will be larger than one. Observing that the values of $u_{1}$ and $u_{2}$ are mainly determined by $k$ and $a$. Recalling that the universe undergoes an exponential expansion during an inflationary period, the small values of $k$ could represent the evolution of Krylov complexity as inflation almost finishes. At the very beginning of inflation, the values of $k$ should be large. Consequently, one could see a complete evolution of Krylov complexity when $k$ varies from tiny values to large values.
@@
First, we will give the precise evolution of Krylov complexity in terms of (60) represented by Fig. 4. One can see that the total trend of Krylov complexity is almost the same with various values of $\alpha$, the only difference is the order of the Krylov complexity, the standard dispersion relation will hold the largest Krylov complexity under the same parameters, which is contradicted with the evolution in closed system presenting in Fig. 3, where the modified case will grow faster compared with the standard case. Observing that there is a blank area for every case, which means that the Krylov complexity will decrease to zero and then it will experience exponential growth. One can also find that the varying trend of Krylov complexity is mainly determined by scale factor $a$, not $k$. In light of this observation, we could fix some specific values of $k$ and vary with respect to $y$ in order to check the evolution trend of Krylov complexity. For completeness, we only consider two cases: (a). $k=0.01$; (b). $k=1000$, which could provide a full perspective of the evolution trend. Fig. 5 clearly indicates that the Krylov complexity will reach the peak of various $\alpha$, then it will dramatically decrease to zero. After some time scale, it will show the exponential growth pattern, in which the standard case will grow faster compared with modified cases. Thus, Fig. 5 will confirm the total varying trend as shown in Fig. 3. In some sense, Fig. 5 has given a precise and complete evolution of Krylov complexity.
Although both the modified and standard cases show an exponential growth trend throughout the inflationary period, it is worth noting that the modified case in the method of open system will grow at a slower pace than the standard case. But the Krylov complexity of the modified case grows faster than that of the standard case in the closed system’s method, as shown in Fig. 3. We need to identify the factors responsible for this difference. The evolution of Fig. 5 comes from the precise formula of Krylov complexity (60). In comparison to the Krylov complexity (37) in the closed system method, it is explicit that the Krylov complexity in open system’s method is influenced by the dissipation effects encoding in dissipative coefficient $u_{2}$. Mathematically speaking, one can easily observe the value of $u_{2}$ (57) is determined by $f^{2}$. If the value of $f^{2}$ is enhanced, the value of Krylov complexity will be suppressed as shown in Eq. (60). Therefore, it is easy to conclude that the strong dissipative effects are responsible for the difference in Krylov complexity’s evolution between the open system method and the closed system method.
@@
Finally, we should compare the Krylov complexity in methods of closed system and open system. The reason we need to investigate the Krylov complexity in a method of open system is that the whole universe is an open system. Ref. Cheung:2007st has shown the universe will break the symmetry along the temporal part, which leads to the non-conservation of energy for the universe. From another perspective, the generation of particle is a non-equilibrium process encoding in the preheating period Kofman:1994rk ; Kofman:1997yn , which is an open system. Based on these two factors, it is quite natural to consider inflation as an open system. Consequently, the method of utilizing the method of an open system for investigating the Krylov complexity is more realistic and reliable compared with the method of closed system.
@@ V.6 Krylov entropy in an open inflationary universe
We have checked that the wave function (58) could nicely recover the case of the weak region, meanwhile it also can restore into the case of a closed system. For completeness, we will also calculate its corresponding Krylov entropy.
Being armed with the definition of Krylov entropy (39) and its resulting formula is derived in Appendix A that reads as follows,
@@
$\begin{split}&S_{k}\approx-\big{(}\cosh^{2}\eta-4u_{2}\cosh^{3}\eta\sinh\eta\big{)}\bigg{[}-2u_{2}\tanh\eta+\rm sech^{2}\eta\ln\rm sech^{2}\eta+2u_{2}\tanh\eta\ln\rm sech^{2}\eta\\ &+\tanh^{2}\eta\ln\tanh^{2}\eta\bigg{]}\\ &=-\sinh^{2}\eta\ln\tanh^{2}\eta-\ln\rm sech^{2}\eta+u_{2}\big{(}2\cosh\eta\sinh\eta+2\cosh\eta\sinh\eta\ln\rm sech^{2}\eta\\ &+4\cosh\eta\sinh^{3}\eta\ln\tanh^{2}\eta\big{)}\\ &=\cosh^{2}\eta\ln\cosh^{2}\eta-\sinh^{2}\eta\ln\sinh^{2}\eta+u_{2}\big{(}2\cosh\eta\sinh\eta+2\cosh\eta\sinh\eta\ln\rm sech^{2}\eta\\ &+4\cosh\eta\sinh^{3}\eta\ln\tanh^{2}\eta\big{)},\end{split}$ (63)
@@
$(a).$ Sec. II gives a brief introduction to Krylov complexity and calculates the Krylov complexity and Lanczos coefficient in the closed system’s method. Sec. III simply introduces the modified dispersion relation in inflation, then we also give some very fundamental information about inflation whose more details can be found in Baumann:2009ds . Based on these two concepts, we found that $b_{n}$ is proportional to $n$ as shown in Eq. (36), thus it could be seen that our universe is an infinite, many-body, and chaotic system confirmed by Bhattacharya:2022gbz . Meanwhile, Fig. 2 indicates that the standard dispersion case of the Lanczos coefficient and Lyapunov index are mainly determined by scale factor (single field inflation), the modified case is mainly determined by momentum $k$, where these two parameters present the chaotic features of the system. To be more precise, the scale factor could mainly determine the growth of the operator and chaos for the standard case. As for the modified case, it was mostly influenced by the momentum.
@@
$(b).$ Fig. 3 clearly indicates that the Krylov complexity will show the irregular oscillation before the horizon exits which shows a similar pattern compared with circuit complexity Li:2023ekd . And the modified case will lead to the faster growth of complexity (exponential growth). We also found that the number density is the same with Krylov complexity similar to Adhikari:2022oxr . Since the number density is proportional to volume, we illustrate the CV conjecture in some sense. Ref. Barbon:2019wsy shows there is a saturation value of Krylov complexity. In our case, it may show a similar trend of complexity by extending our analysis into RD and MD. Meanwhile, the method what we consider is for a closed system.
@@
$(c).$ Refs. Cheung:2007st ; Kofman:1994rk ; Kofman:1997yn has indicated that our universe is an open sytem. Thus, the method of open system treating the Krylov complexity is more realistic and reliable compared with that of closed system. In the approach of an open system, the investigation is based on our central result, an exact wave function (58) describing the strong and weak dissipative system, including inflation. Compared with the original method for constructing the wave function Bhattacharjee:2022lzy , our wave function (58) could describe the strong and weak dissipative region. According to $\eta=-\frac{1}{aH}$, one can see that $\eta$ cannot be larger than zero in inflation. But the situation may be different after inflation since $\eta\propto t^{\alpha}$ in the late universe. Thus, one can conclude that the inflation is a strong dissipative system. However, the analysis can be dubbed as a criterion for testing the validity of the exact wave function (58). As a further test, the calculation for Krylov complex (61) and Krylov entropy (63) could nicely recover into the case of a closed system. In the weak dissipative region, our calculation could return into the analysis of Bhattacharjee:2022lzy , where there is scrambling time dubbed as a boundary for studying the complexity. The scrambling time is just $\xi$ (54) which is a variable determined by $k$ and $a$. For the strong dissipative region, one could see that the varying trend is always growing from the large scales, which will not reach some constant values. Fig. 4 clearly indicates the evolution trend of Krylov complexity. Generally, one cannot distinguish the standard case and modified case from the strong dissipative region. At the small scale, there is also one peak of Krylov complexity. One can see that inflation will have a huge impact on the Krylov complexity.
@@
Secondly, all of our analysis is based on the Hamilton (23) that is Hermitian. The original method for evaluating the Krylov complexity in an open system is based on the Lindblad master equation Lindblad:1975ef ; Gorini:1975nb . The corresponding Lindblad of various dynamical systems is non-Hermitian. For obtaining the non-Hermitian Hamilton, one could expand the action into the higher order in terms of the Mukhanov variable. Another way is that we could add the potential of inflaton in MD or RD since the contribution of potential cannot be ignored due to the violation of slow-roll condition.
Thirdly, the horizon exits will lead to the curvature perturbation transit from the quantum level to the classical level. Thus, this process may impact the evolution of Krylov complexity. Especially, one can expand the action of inflaton to the higher order to produce the minimal decoherence Burgess:2022nwu , which only incorporates the self-interaction via General Relativity. We are still curious about whether the holographic principle can be applied to this inflationary period or not, just like the CV or CA conjecture.
@@ Appendix A The calculation of Krylov entropy
$\begin{split}&S_{k}=-\frac{1}{A^{2}}\bigg{[}\rm sech^{2}\eta[1+2u_{2}\tanh\eta+(u_{2}^{2}-|1-u_{1}^{2}|)\tanh^{2}\eta][\ln\rm sech^{2}-\ln(1+u_{2}\tanh\eta)^{2}]\\ &+\rm sech^{2}\eta\tanh^{2}\eta|1-u_{1}^{2}|\ln|1-u_{1}^{2}|+|1-u_{1}^{2}|\tanh^{2}\eta\rm sech^{2}\eta[\ln\tanh^{2}\eta-\ln(1+u_{2}\tanh\eta)^{2}]\bigg{]}\\ &=-\frac{1}{A^{2}}\bigg{[}-\ln(1+u_{2}\tanh\eta)^{2}(\rm sech^{2}\eta+2u_{2}\rm sech^{2}\eta\tanh\eta+u_{2}^{2}\rm sech^{2}\eta\tanh^{2}\eta)\\ &+\rm sech^{2}\eta\ln\rm sech^{2}\eta[1+2u_{2}\tanh\eta+(u_{2}^{2}-|1-u_{1}^{2}|)\tanh^{2}\eta]+\rm sech^{2}\eta\tanh^{2}\eta|1-u_{1}^{2}|\ln|1-u_{1}^{2}|\\ &+|1-u_{1}^{2}|\tanh^{2}\eta\rm sech^{2}\eta\ln\tanh^{2}\eta\bigg{]}\\ &=\frac{\rm sech^{2}\eta}{A^{2}}\bigg{[}(1+u_{2}\tanh\eta)^{2}\ln(1+u_{2}\tanh\eta)^{2}-[(1+u_{2}\tanh\eta)^{2}-|1-u_{1}^{2}|\tanh^{2}\eta]\ln\rm sech^{2}\eta\\ &-\tanh^{2}\eta|1-u_{1}^{2}|\ln|1-u_{1}^{2}|-|1-u_{1}^{2}|\tanh^{2}\eta\ln\tanh^{2}\eta\bigg{]}.\end{split}$ (70)